# SFTMix: Elevating Language Model Instruction Tuning with Mixup Recipe

### v2

@@ Abstract
To acquire instruction-following capabilities, large language models (LLMs) undergo instruction tuning, where they are trained on instruction-response pairs using next-token prediction (NTP). Efforts to improve instruction tuning often focus on higher-quality supervised fine-tuning (SFT) datasets, typically requiring data filtering with proprietary LLMs or human annotation. In this paper, we take a different approach by proposing SFTMix, a novel Mixup-based recipe that elevates LLM instruction tuning beyond the conventional NTP paradigm, without relying on well-curated datasets. Observing that LLMs exhibit uneven confidence across the semantic representation space, we argue that examples with different confidence levels should play distinct roles in instruction tuning—confident data is prone to overfitting, while unconfident data is harder to generalize. Based on this insight, SFTMix leverages training dynamics to identify examples with varying confidence levels, interpolates them to bridge the confidence gap, and applies a Mixup-based regularization to support learning on these additional, interpolated examples. By propagating supervision signals across confidence regions and encouraging linear behavior between them, SFTMix mitigates overfitting in confident examples while enhancing generalization in unconfident ones. We demonstrate the effectiveness of SFTMix in both instruction-following and healthcare-specific SFT tasks, with consistent improvements across LLM families and SFT datasets of varying sizes and qualities. Extensive analyses across six directions highlight SFTMix’s compatibility with data selection, adaptability to compute-constrained scenarios, and scalability to broader applications.
@@
Yuxin Xiao1thanks: Work done during an internship at Zoom., Shujian Zhang2, Wenxuan Zhou2, Marzyeh Ghassemi1, Sanqiang Zhao2thanks: Corresponding author. 1Massachusetts Institute of Technology, 2Zoom Video Communications {yuxin102, mghassem}@mit.edu {shujian.zhang, wenxuan.zhou, sanqiang.zhao}@zoom.com
@@ 1 Introduction
Large language models (LLMs) have recently achieved remarkable performance across a wide range of natural language processing (NLP) tasks (Zhao et al., 2023; Minaee et al., 2024). After being pre-trained on large corpora of raw text, LLMs undergo a critical instruction-tuning stage (Ouyang et al., 2022; Zhang et al., 2023) to develop their instruction-following capabilities based on supervised fine-tuning (SFT) datasets (Taori et al., 2023; Wang et al., 2023; Xu et al., 2024). During this stage, LLMs are usually trained via next-token prediction (NTP), where they predict the next token in a response given both the instruction and the preceding tokens in that response.
@@
Previous research in this field has predominantly focused on enhancing the quality of instruction-tuning datasets. One line of research direction seeks to better understand the intrinsic properties of these datasets (Kung et al., 2023; Lin et al., 2024) and selects informative instruction-response pairs through heuristic-based filters (Zhao et al., 2024) or LLM scoring (Chen et al., 2024). Another line of work generates high-quality responses by querying advanced proprietary LLMs (Chen et al., 2024) or relying on human annotators (Zhou et al., 2023).
@@
In this paper, we take a different approach by proposing SFTMix, a novel Mixup-based (Zhang et al., 2018) recipe to elevate LLM instruction tuning beyond the conventional NTP paradigm, without the need for well-curated datasets. Our design is motivated by the observation that an LLM’s confidence distribution is uneven across the semantic representation space. Since confident data is prone to overfitting (Zhang and Vaidya, 2021; Han et al., 2024) and unconfident data is harder to generalize (Elsayed et al., 2018; Jiang et al., 2018), we argue that data with varying confidence levels should play distinct roles in instruction tuning. Hence, we first derive an LLM’s confidence from its training dynamics (Swayamdipta et al., 2020) and divide the SFT dataset into confident and unconfident subsets accordingly. We then linearly interpolate between these subsets and introduce a Mixup-based regularization to support learning on these additional, interpolated examples. By propagating supervision signals across confidence regions (Bengio et al., 2009; Chapelle et al., 2009; Sohn et al., 2020) and encouraging linear behavior between them (Zhang et al., 2018; Verma et al., 2019), our recipe mitigates overfitting in confident examples while enhancing generalization in unconfident ones during LLM instruction tuning.
@@
We demonstrate the effectiveness of our proposed SFTMix recipe in both instruction-following and domain-specific SFT settings. In particular, SFTMix significantly outperforms the conventional NTP baseline in both MT-Bench (Zheng et al., 2023) and AlpacaEval-2 (Dubois et al., 2024), with consistent improvements across LLM families (i.e., Llama (Dubey et al., 2024), Mistral (Jiang et al., 2023)) and SFT datasets of varying sizes and qualities (i.e., Alpaca-52K (Taori et al., 2023), UltraChat-200K (Tunstall et al., 2023)). Moreover, in the healthcare domain, Llama-3.1-8B and Mistral-7B-v0.1, instruction-tuned on MedAlpaca-263K (Han et al., 2023) using SFTMix, achieve an average of $1.5\%$ absolute increase in accuracy across four question-answering benchmarks (Jin et al., 2019, 2021; Pal et al., 2022).
@@
In addition, we conduct in-depth analyses across six directions to highlight SFTMix’s versatility and scalability in LLM instruction tuning. Our findings validate the importance of confidence-based data splitting for effective Mixup and demonstrate that Mixup works best as a regularization alongside NTP. Furthermore, we show that SFTMix integrates seamlessly with data selection methods, adapts well to compute-constrained scenarios, and scales effectively to broader applications.
@@
  • We show that SFTMix outperforms the NTP baseline across various instruction-following and healthcare-specific SFT tasks, with consistent improvements across LLM families and SFT datasets of varying size and quality.
  • Extensive analyses across six directions highlight SFTMix’s compatibility with data selection, adaptability to compute-constrained scenarios, and scalability to broader applications.
@@ 2 Related Work
To align LLMs with user intents or domain-specific tasks, Ouyang et al. (2022) propose instruction-tuning LLMs on human-annotated demonstrations using supervised learning. The conventional NTP paradigm trains LLMs to predict response tokens sequentially given instruction-response pairs (Zhang et al., 2023). Enhancements include adding noise to token embeddings (Jain et al., 2024), commonality-aware partition (Rao et al., 2024), and explicitly modeling instructions (Shi et al., 2024). Previous work (Chiang et al., 2023; Ding et al., 2023; Taori et al., 2023; Wang et al., 2023; Xu et al., 2024) collects instruction-following datasets via LLM distillation or crowd-sourced user conversations. To improve data quality, researchers employ heuristic-based filters (Schoch et al., 2023; Zhao et al., 2024), importance weighting (Xie et al., 2023; Xia et al., 2024), LLM scoring (Chen et al., 2024), and human curation (Zhou et al., 2023). Other studies explore the intrinsic properties of SFT datasets (Kung et al., 2023; Lin et al., 2024). However, acquiring high-quality SFT data often entails substantial computational and labor costs. This paper aims to optimize data utilization through insightful data interpretation and elevate instruction tuning beyond NTP without relying on well-curated datasets.
@@
Data characterization (Albalak et al., 2024; Wang et al., 2024) analyzes training data quality to improve downstream model performance. In particular, Swayamdipta et al. (2020) leverage training dynamics from a pre-trained language model (Liu et al., 2019) to create data maps, inspiring advances in active learning (Zhang and Plank, 2021; Zhang et al., 2022; Kung et al., 2023), curriculum learning (Christopoulou et al., 2022; Lin et al., 2024; Poesina et al., 2024), and dataset pruning (Chimoto et al., 2024; He et al., 2024; Lin et al., 2024; Seedat et al., 2024). Here, we apply training dynamics to causal language generation by categorizing an SFT dataset into confident and unconfident subsets, which facilitates the subsequent Mixup-based regularization during LLM instruction tuning.
@@
To mitigate memorization and adversarial sensitivity, Zhang et al. (2018) propose Mixup, which trains models on convex combinations of pairs of input features and their corresponding labels. Its variants (Verma et al., 2019; Hendrycks et al., 2020; Uddin et al., 2021; Choi et al., 2022) further suggest interpolating feature representations at different stages, guided by various training signals. Theoretical analyses (Zhang et al., 2021; Carratino et al., 2022; Chidambaram et al., 2022; Park et al., 2022; Pinto et al., 2022) highlight its data-adaptive regularization and generalization effects, leading to strong out-of-distribution robustness and well-calibrated uncertainty estimation. Empirical studies confirm its effectiveness in semi-supervised learning (Berthelot et al., 2019, 2020; Li et al., 2020, 2022) and NLP (Chen et al., 2020; Guo et al., 2020; Sun et al., 2020; Park and Caragea, 2022; Yang et al., 2022). We extend Mixup to LLM instruction tuning, proposing a regularization method to reduce overfitting to confident examples while supporting learning for less confident ones.
@@ 3 SFTMix
Based on the preliminaries in §3.1, we discuss the motivation in §3.2 and introduce SFTMix in §3.3.
@@ 3.1 Preliminaries
Figure 1: Embeddings of $2{,}500$ most and $2{,}500$ least confident examples in Alpaca-52K by Llama-3.1-8B trained using NTP. The clear separation between these embeddings suggests that the LLM exhibits varying confidence levels across different semantic regions.
@@
3.2 Motivation
We motivate the design of SFTMix through a case study. Specifically, we instruction-tune Llama-3.1-8B (Dubey et al., 2024) on Alpaca-52K (Taori et al., 2023) and collect the LLM’s confidence for each training data point across five checkpoints. Using the last hidden state $\mathbf{Z}$ of the final token in $(\mathcal{X}_{i},\mathcal{Y}_{i})$ as its semantic representation, we visualize $2{,}500$ most and $2{,}500$ least confident examples via t-SNE (Van der Maaten and Hinton, 2008). As shown in Figure 1, embeddings of data points with contrasting confidence levels are clearly separated, indicating that the LLM exhibits uneven confidence across the semantic representation space.
Furthermore, we analyze the distributions of instruction topics in the $50$ most and $50$ least confident examples. We find that the most confident examples primarily involve deterministic grammar tasks (e.g., "correct any grammar error in the following sentence"), while $56\%$ of the least confident examples require creative content generation (e.g., "find a name for an e-commerce website"), and the remaining $44\%$ consist of noisy or unanswered instructions. This aligns with our observation in Figure 1, showing that the LLM’s confidence varies across different semantic regions.
@@ 3.3 Recipe
Here, $\mu$ is a hyperparameter to control the regularization effect. Since $\nabla_{\mathbf{W}}H(\tilde{\mathbf{Y}}_{n},\sigma(\tilde{\mathbf{Z}}_{n}^{\top}\mathbf{W}))=\tilde{\mathbf{Z}}_{n}^{\top}(\sigma(\tilde{\mathbf{Z}}_{n}^{\top}\mathbf{W})-\tilde{\mathbf{Y}}_{n})$ involves the nonlinear softmax operation $\sigma$, $\ell_{\text{Mixup}}$ modifies the gradient descent direction in NTP by incorporating the interpolated examples (see §A for the derivation details). For ease of implementation, we ensure that each training batch contains an equal number of confident and unconfident examples, which are then randomly paired for the linear interpolation and the Mixup-based regularization.
@@ 4.1 Instruction-Following SFT
We compare SFTMix with NTP by applying them to the instruction tuning of two pre-trained LLMs (i.e., Llama-3.1-8B (Llama) (Dubey et al., 2024) and Mistral-7B-v0.1 (Mistral) (Jiang et al., 2023)) on two instruction-following datasets of varying scales and qualities (i.e., the uncurated Alpaca-52K (Taori et al., 2023) and the filtered UltraChat-200K (Tunstall et al., 2023)). We then evaluate the instruction-tuned LLMs on two instruction-following benchmarks: MT-Bench (Zheng et al., 2023) and AlpacaEval-2 (Dubois et al., 2024). Following Zhao et al. (2024), we also conduct a human evaluation for head-to-head comparisons using the Vicuna subset in AlpacaEval-2.
Implementation Details.
By default, we use a separate instance of the same type as the target instruction-tuning LLM to obtain $\text{Conf}\,(\mathcal{Y}_{i}\,|\,\mathcal{X}_{i})$ in Step 1 of SFTMix. We train each LLM on Alpaca-52K for three epochs and on UltraChat-200K for one epoch, using a batch size of $32$ on eight H100 GPUs. The instruction-tuning process leverages the AdamW optimizer with a learning rate of $2\mathrm{e}{-6}$ and a weight decay of $0.1$. We also apply a cosine learning rate scheduler with a warm-up ratio of $0.1$. Based on our hyperparameter search in §B.1, we set $\alpha=0.5$ for sampling $\lambda$ and $\mu=0.2$ when constructing $\ell_{\text{SFTMix}}$. The NTP baseline follows the same setup but excludes the Mixup-based regularization $\ell_{\text{Mixup}}$. When training on UltraChat-200K, we expand each multi-turn interaction into multiple single-turn interactions by incorporating the chat history into the instructions. In MT-Bench and AlpacaEval-2, we employ GPT-4 (Achiam et al., 2023) for LLM-as-a-judge and report the results averaged over five evaluation rounds.
@@
Table 1: Evaluation of instruction-following capabilities of LLMs trained with NTP or SFTMix. We report the average score over five evaluation rounds, highlighting the best-performing instruction-tuning recipe in bold. SFTMix outperforms NTP on MT-Bench and AlpacaEval-2 across both instruction-tuning datasets and LLMs.
@@
As illustrated in Table 1, instruction-tuning with SFTMix consistently outperforms NTP across all metrics in both evaluation benchmarks, regardless of the base LLM or SFT dataset. Notably, SFTMix yields a greater improvement in the multi-turn (MT) conversational ability (an average increase of $0.3$) compared to single-turn (ST) performance (an average increase of $0.2$) in MT-Bench. In AlpacaEval-2, the improvement is particularly evident in the length-controlled win rate (LC WR), which better aligns with human judgment by adjusting for GPT-4’s preference for longer responses. While instruction-tuning with the larger, higher-quality UltraChat-200K dataset results in higher overall scores in MT-Bench and raw win rates (WRs) in AlpacaEval-2, it also produces longer responses, leading to relatively lower LC WRs. Moreover, our human evaluation indicates that instruction-tuning with SFTMix wins $42.5\%$ of the head-to-head comparisons, while NTP wins only $26.5\%$ (details in §B.2). This agrees with the conclusion from LLM-as-a-judge evaluations.
@@
[…]

### v3

@@ Abstract
To acquire instruction-following capabilities, large language models (LLMs) undergo instruction tuning, where they are trained on instruction-response pairs using next-token prediction (NTP). Efforts to improve instruction tuning often focus on higher-quality supervised fine-tuning (SFT) datasets, typically requiring data filtering with proprietary LLMs or human annotation. In this paper, we take a different approach by proposing SFTMix, a novel Mixup-based recipe that elevates LLM instruction tuning without relying on well-curated datasets. We observe that LLMs exhibit uneven confidence across the semantic representation space. We argue that examples with different confidence levels should play distinct roles in instruction tuning: Confident data is prone to overfitting, while unconfident data is harder to generalize. Based on this insight, SFTMix leverages training dynamics to identify examples with varying confidence levels. We then interpolate them to bridge the confidence gap and apply a Mixup-based regularization to support learning on these additional, interpolated examples. We demonstrate the effectiveness of SFTMix in both instruction-following and healthcare-specific SFT tasks, with consistent improvements across LLM families and SFT datasets of varying sizes and qualities. Extensive analyses across six directions highlight SFTMix’s compatibility with data selection, adaptability to compute-constrained scenarios, and scalability to broader applications.
@@
Yuxin Xiao1thanks: Work done during an internship at Zoom., Shujian Zhang2, Wenxuan Zhou2, Marzyeh Ghassemi1, Sanqiang Zhao2thanks: Corresponding author. 1Massachusetts Institute of Technology, 2Zoom Video Communications yuxin102@mit.edu
@@ 1 Introduction
LLMs have recently achieved strong performance across diverse natural language processing (NLP) tasks (Zhao et al., 2023; Minaee et al., 2024). After being pre-trained on large corpora of raw text, LLMs undergo a critical instruction-tuning stage (Ouyang et al., 2022; Zhang et al., 2023) to develop their instruction-following capabilities based on SFT datasets (Taori et al., 2023; Wang et al., 2023; Xu et al., 2024). During this stage, LLMs are usually trained via NTP, where they predict the next token in a response given both the instruction and the preceding tokens in that response.
@@
Previous research in this field has predominantly focused on enhancing the quality of instruction-tuning datasets. One line of research direction seeks to better understand the intrinsic properties of these datasets (Kung et al., 2023; Lin et al., 2024) and selects informative instruction-response pairs through heuristic-based filters (Zhao et al., 2024) or LLM scoring (Chen et al., 2024a). Another line of work generates high-quality responses by querying advanced proprietary LLMs (Chen et al., 2024a) or relying on human annotators (Zhou et al., 2023).
@@
In this paper, we take a different approach by proposing SFTMix, a novel Mixup-based (Zhang et al., 2018) recipe to elevate LLM instruction tuning without the need for well-curated datasets. Our design is motivated by the observation that an LLM’s confidence distribution is uneven across the semantic representation space. Since confident data is prone to overfitting (Zhang and Vaidya, 2021; Han et al., 2024) and unconfident data is harder to generalize (Elsayed et al., 2018; Jiang et al., 2018), we argue that data with varying confidence levels should play distinct roles in instruction tuning. Hence, we first derive an LLM’s confidence from its training dynamics (Swayamdipta et al., 2020) and divide the SFT dataset into confident and unconfident subsets accordingly. We then linearly interpolate between these subsets and introduce a Mixup-based regularization to support learning on these additional, interpolated examples. By propagating supervision signals across confidence regions (Bengio et al., 2009; Chapelle et al., 2009; Sohn et al., 2020) and encouraging linear behavior between them (Zhang et al., 2018; Verma et al., 2019), our recipe mitigates overfitting in confident examples while enhancing generalization in unconfident ones during LLM instruction tuning.
@@
We demonstrate the effectiveness of our proposed SFTMix recipe in both instruction-following and domain-specific SFT settings. In particular, SFTMix significantly outperforms the conventional NTP baseline in both MT-Bench (Zheng et al., 2023) and AlpacaEval-2 (Dubois et al., 2024), with consistent improvements across LLM families (i.e., Llama (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Qwen (Hui et al., 2024)) and SFT datasets of varying sizes and qualities (i.e., Alpaca-52K (Taori et al., 2023), UltraChat-200K (Tunstall et al., 2023), and Tulu3-939K (Lambert et al., 2024)). Moreover, in the healthcare domain, Llama-3.1-8B and Mistral-7B-v0.1, instruction-tuned on MedAlpaca-263K (Han et al., 2023) using SFTMix, increase the accuracy by an average of $1.5\%$ absolutely across four question-answering benchmarks (Jin et al., 2019, 2021; Pal et al., 2022).
@@
In addition, we conduct in-depth analyses across six directions to highlight SFTMix’s versatility and scalability in LLM instruction tuning. Our results validate the importance of confidence-based data splitting for effective Mixup and show that Mixup works best as a regularization alongside NTP. Moreover, we demonstrate that SFTMix integrates seamlessly with data selection (Chen et al., 2024a; Zhao et al., 2024), adapts well to compute-constrained scenarios (Hu et al., 2022), and scales effectively to broader applications (Burns et al., 2024).
@@
  • We show that SFTMix outperforms the NTP baseline across various instruction-following and healthcare-specific SFT tasks, with consistent improvements across LLM families and SFT datasets of varying sizes and qualities.
  • Extensive analyses across six directions highlight that SFTMix is compatible with data selection, adaptable to compute-constrained scenarios, and scalable to broader applications.
@@ 2 Related Work
To align LLMs with user intents or domain-specific tasks, Ouyang et al. (2022) propose instruction-tuning LLMs on human-annotated demonstrations using supervised learning. The conventional NTP paradigm trains LLMs to predict response tokens sequentially given instruction-response pairs (Zhang et al., 2023). Enhancements include adding noise to token embeddings (Jain et al., 2024), commonality-aware partition (Rao et al., 2024), and explicitly modeling instructions (Shi et al., 2024). Previous work (Chiang et al., 2023; Ding et al., 2023; Taori et al., 2023; Wang et al., 2023; Xu et al., 2024) collects instruction-following datasets via LLM distillation or crowd-sourced user conversations. To improve data quality, researchers employ heuristic-based filters (Schoch et al., 2023; Zhao et al., 2024; Yang et al., 2025), importance weighting (Xie et al., 2023; Xia et al., 2024; Chen et al., 2024b), LLM scoring (Chen et al., 2024a), and human curation (Zhou et al., 2023). Other studies explore the intrinsic properties of SFT datasets (Kung et al., 2023; Lin et al., 2024; Fu et al., 2025; Zhang et al., 2025a), gradient-based methods (Wang et al., 2025; Zhao et al., 2025), and reinforcement learning from human feedback (Rafailov et al., 2023; Ethayarajh et al., 2024; Zeng et al., 2024). However, acquiring high-quality SFT data often entails substantial computational and labor costs. This paper aims to optimize data utilization through confidence-aware data interpretation and elevate instruction tuning beyond NTP without relying on well-curated datasets.
@@
Data characterization (Albalak et al., 2024; Wang et al., 2024) analyzes training data quality to improve downstream model performance. In particular, Swayamdipta et al. (2020) leverage training dynamics from a pre-trained language model (Liu et al., 2019) to create data maps. This idea has inspired advances in active learning (Zhang and Plank, 2021; Zhang et al., 2022; Kung et al., 2023), curriculum learning (Christopoulou et al., 2022; Lin et al., 2024; Poesina et al., 2024), dataset pruning (Chimoto et al., 2024; He et al., 2024; Lin et al., 2024; Seedat et al., 2024), and LLM scaling (Mircea et al., 2025; Qi et al., 2025; Zhang et al., 2025b). Here, we apply training dynamics to causal language generation by categorizing an SFT dataset into confident and unconfident subsets, which facilitates the subsequent Mixup-based regularization during LLM instruction tuning.
@@
To mitigate memorization and adversarial sensitivity, Mixup Zhang et al. (2018) trains models on convex combinations of paired inputs and labels. Its variants (Verma et al., 2019; Hendrycks et al., 2020; Uddin et al., 2021; Choi et al., 2022) interpolate feature representations at various stages, guided by different training signals. Theoretical analyses (Zhang et al., 2021; Carratino et al., 2022; Chidambaram et al., 2022; Park et al., 2022; Pinto et al., 2022) highlight its adaptive regularization and generalization effects, yielding stronger out-of-distribution robustness and better uncertainty calibration. Empirical studies confirm its effectiveness in semi-supervised learning (Berthelot et al., 2019, 2020; Li et al., 2020, 2022) and NLP (Chen et al., 2020; Guo et al., 2020; Sun et al., 2020; Park and Caragea, 2022; Yang et al., 2022). In this paper, we extend Mixup to LLM instruction tuning, proposing a regularization method to reduce overfitting to confident examples while supporting learning for less confident ones.
@@ 3 SFTMix
Based on the preliminaries in §3.1, we discuss the motivation in §3.2, introduce SFTMix in §3.3, and analyze its effect on the tuning process in §3.4.
@@ 3.1 Preliminaries
Figure 1: Embeddings of 2,5002{,}500 most and 2,5002{,}500 least confident examples in Alpaca-52K by Llama-3.1-8B trained using NTP. The clear separation between these embeddings suggests that the LLM exhibits varying confidence levels across different semantic regions.
3.2 Motivation
We motivate the design of SFTMix through a case study. Specifically, we instruction-tune Llama-3.1-8B (Dubey et al., 2024) on Alpaca-52K (Taori et al., 2023) and collect the LLM’s confidence for each training data point across five checkpoints. Using the last hidden state $\mathbf{Z}$ of the final token in $(\mathcal{X}_{i},\mathcal{Y}_{i})$ as its semantic representation, we visualize 2,5002{,}500 most and 2,5002{,}500 least confident examples via t-SNE (Van der Maaten and Hinton, 2008). As shown in Figure 1, embeddings of data points with contrasting confidence levels are clearly separated, indicating that the LLM exhibits uneven confidence across the semantic representation space.
@@ 3.2 Motivation
Furthermore, we analyze the distributions of instruction topics in the $50$ most and $50$ least confident examples. We find that the most confident examples primarily involve deterministic grammar tasks (e.g., “correct any grammar error in the following sentence”), while $56\%$ of the least confident examples require creative content generation (e.g., “find a name for an e-commerce website”), and the remaining $44\%$ consist of noisy or unanswered instructions. This aligns with our observation in Figure 1, showing that the LLM’s confidence varies across different semantic regions.
@@ 3.3 Recipe
Here, $\mu$ is a hyperparameter to control the regularization effect. For ease of implementation, we ensure that each training batch contains an equal number of confident and unconfident examples, which are then randomly paired for the linear interpolation and the Mixup-based regularization.
3.4 Analysis
Here, we analyze how the Mixup-based regularization affects the direction of gradient descent in the original NTP paradigm. For notational simplicity, we focus on the cross-entropy $H(\tilde{\mathbf{Y}},\sigma(\tilde{\mathbf{Z}}^{\top}\mathbf{W}))$ between the interpolated one-hot encoding vector $\tilde{\mathbf{Y}}$ and the corresponding interpolated representation $\tilde{\mathbf{Z}}$ from the target LLM’s last transformer layer.
Let $\tilde{\mathbf{S}}=\tilde{\mathbf{Z}}^{\top}\mathbf{W}$, the gradient of $H$ w.r.t. $\tilde{\mathbf{S}}$ is
Using the chain rule, we have
[…]
In other words, the gradient from the regularization with the interpolated example does not decompose into a weighted sum of the gradients from the NTP loss on the corresponding confident and unconfident examples. As a result, the Mixup-based regularization modifies the gradient descent direction in NTP by incorporating these interpolated examples.
@@ 4.1 Instruction-Following SFT
We compare SFTMix with NTP and a sample-reweighting baseline, IR-DRO (Chen et al., 2024b), using three pre-trained LLMs: Llama-3.1-8B (Llama) (Dubey et al., 2024), Mistral-7B-v0.1 (Mistral) (Jiang et al., 2023), and Qwen-2.5-14B (Qwen) (Hui et al., 2024). Due to computational constraints, we train Qwen only on the smaller, uncurated Alpaca-52K (Taori et al., 2023), while Llama and Mistral are trained on both Alpaca-52K and the larger filtered datasets: UltraChat-200K (Tunstall et al., 2023) and Tulu3-939K (Lambert et al., 2024). For the same reason, IR-DRO is evaluated only on Llama and Mistral using Alpaca-52K. We then evaluate the instruction-tuned LLMs on two instruction-following benchmarks: MT-Bench (Zheng et al., 2023) and AlpacaEval-2 (Dubois et al., 2024). Following Zhao et al. (2024), we also conduct a human evaluation for head-to-head comparisons using the Vicuna subset in AlpacaEval-2.
@@
Dataset: Tulu3-939K
Llama NTP $6.2500$ $4.9625$ $5.6063$ $7.1045$ $12.0811$
SFTMix 6.6721 5.3331 6.0026 8.0129 13.1345
Mistral NTP $5.9246$ $4.0128$ $4.9687$ $6.9926$ $11.2471$
SFTMix 6.3626 4.3398 5.3512 7.5421 11.8681
Table 1: Evaluation of instruction-following capabilities of LLMs trained with NTP, SFTMix, and IR-DRO. We highlight the best-performing instruction-tuning recipe in bold. SFTMix outperforms the baselines consistently across instruction-tuning datasets and LLMs.
Implementation Details.
By default, we use a separate instance of the same type as the target instruction-tuning LLM to obtain $\text{Conf}\,(\mathcal{Y}_{i}\,|\,\mathcal{X}_{i})$ in Step 1 of SFTMix. We train each LLM on Alpaca-52K for three epochs and on UltraChat-200K and Tulu3-939K for one epoch, with a batch size of $32$ on eight H100 GPUs. The tuning process employs the AdamW optimizer with a learning rate of $2\mathrm{e}{-6}$, a weight decay of $0.1$, and a cosine learning rate scheduler with a warm-up ratio of $0.1$. Based on our hyperparameter search in §A.1, we set $\alpha=0.5$ for sampling $\lambda$ and $\mu=0.2$ when constructing $\ell_{\text{SFTMix}}$. The NTP baseline follows the same setup but excludes the Mixup-based regularization $\ell_{\text{Mixup}}$. When training on UltraChat-200K and Tulu3-939K, we expand each multi-turn interaction into multiple single-turn interactions by incorporating the chat history into the instructions. In MT-Bench and AlpacaEval-2, we employ GPT-4 (Achiam et al., 2023) for LLM-as-a-judge and report the results averaged over five evaluation rounds.
@@
As illustrated in Table 1, instruction-tuning with SFTMix consistently outperforms NTP and IR-DRO across all metrics in both evaluation benchmarks, regardless of the base LLM or SFT dataset. Notably, SFTMix yields a greater improvement in the multi-turn (MT) conversational ability (an average increase of $0.32$) compared to single-turn (ST) performance (an average increase of $0.27$) in MT-Bench. In AlpacaEval-2, the improvement is particularly evident in the length-controlled win rate (Dubois et al., 2024) (LC WR), which better aligns with human judgment by adjusting for GPT-4’s preference for longer responses. While instruction-tuning with the larger, higher-quality UltraChat-200K dataset results in higher scores in MT-Bench and raw win rates (WRs) in AlpacaEval-2, it also produces longer responses, leading to relatively lower LC WRs. In contrast, SFTMix with Tulu3-939K delivers larger gains across all metrics than UltraChat-200K.
Further Analysis.
Our human evaluation indicates that instruction-tuning with SFTMix wins $42.5\%$ of the head-to-head comparisons, while NTP wins only $26.5\%$ (details in §A.2). This agrees with the conclusion from LLM-as-a-judge evaluations. We also provide qualitative examples from MT-Bench in §A.3 to illustrate these differences and demonstrate the effectiveness of SFTMix in multilingual instruction tuning in §A.4. Furthermore, we compare the confidence distribution of Llama instruction-tuned with SFTMix versus NTP on Alpaca-52K. SFTMix reduces the standard deviation of confidence scores by $7\%$, indicating a more uniform distribution. This suggests SFTMix helps prevent overfitting on confident samples while better supporting learning from less confident ones.
@@
Table 2: Evaluation results on four healthcare-related benchmarks by prior biomedical LLMs and LLMs trained using either NTP or SFTMix. We bold the scores from the best-performing instruction-tuning recipe. SFTMix achieves a $1.5\%$ absolute increase in average accuracy compared to NTP for both Llama and Mistral.